OpenStudy (anonymous):

how would you begin to sketch the graph of the equation. x = 3y2 − 9

OpenStudy (campbell_st):

well I would look at the fact the x-axis is the line of symmetry factor it x=(y -3)(y + 3) so now I know the y- intercepts... and when y = 0 the x intercept is x = -9 I also know the curve is concave right hope it helps

OpenStudy (triciaal):

one approach put in the form y = find the value of y when x = 0 point (0. y) find the value of x when y = 0 point (x, 0) need the minimum of 2 points to sketch a line you will now have these 2 points

OpenStudy (campbell_st):

ooops... for got yo say its a parabola

OpenStudy (triciaal):

@campbell_st since the x and y 's are switched would this be a horizontal parabola?

OpenStudy (campbell_st):

opps made a mistake with the y intercepts... misread the question the factored from is \[x = 3(y - \sqrt{3})(y + \sqrt{3})\]

OpenStudy (campbell_st):

it is a parabola that opens to the right... and easy way to look at it is with https://www.desmos.com/calculator

OpenStudy (triciaal):

@campbell_st thank you